# Engagement, User Satisfaction, and the Amplification of Divisive Content on Social Media

### v5

@@
Smitha Milli Cornell Tech Corresponding author: smilli@cornell.edu Micah Carroll University of California, Berkeley Yike Wang University of California, Berkeley
@@
University of California, Berkeley
Sebastian Zhao University of California, Berkeley Anca D. Dragan University of California, Berkeley
@@
In a pre-registered randomized experiment, we found that, relative to a reverse-chronological baseline, Twitter’s engagement-based ranking algorithm amplifies emotionally charged, out-group hostile content that users say makes them feel worse about their political out-group. Furthermore, we find that users do not prefer the political tweets selected by the algorithm, suggesting that the engagement-based algorithm underperforms in satisfying users’ stated preferences. Finally, we explore the implications of an alternative approach that ranks content based on users’ stated preferences and find a reduction in angry, partisan, and out-group hostile content but also a potential reinforcement of echo chambers. The evidence underscores the necessity for a more nuanced approach to content ranking that balances engagement, users’ stated preferences, and sociopolitical outcomes.
Table 1: Prior and Current Research on Algorithmic Ranking Outcomes
Ranking by engagement Ranking by stated pref.
No. Outcome Obs. Experiment Experiment
Sock puppets Users Users
1 Emotions in content  (?, ?, ?, ?, ?, ?)  (?) Ours, (?) Ours
2 Partisanship of content  (?, ?, ?, ?, ?, ?)  (?, ?) Ours, (?, ?, ?, ?) Ours
3 Out-group animosity in content  (?, ?) Ours Ours
4 Reader’s emotions  (?, ?) Ours Ours
5 Reader’s in-group perception  (?) Ours, (?) Ours
6 Reader’s out-group perception  (?) Ours, (?, ?) Ours
7 Reader’s stated pref. for content Ours, (?) Ours, (?, ?)
The outcomes analyzed in our paper and prior work. Prior research mainly studied engagement-based ranking, with both observational studies and controlled experiments that used either sock puppets or real users. To our knowledge, we are the first to study the implications of ranking by users’ stated preferences on these outcomes (and thus, for brevity, we omit the empty columns for sock puppet experiments and observational studies). We classified a study as providing observational evidence for the effects of engagement-based ranking if it investigated an association between the amount of engagement content received and any of the listed outcomes. Black cells indicate instances where we could not find any relevant prior work.
Social media ranking algorithms personalize content to individual user preferences. However, many hypothesize that these algorithms could inadvertently facilitate the propagation of angry content, stoke hostile discourse, and escalate polarization (?, ?, ?, ?). Although numerous studies have delved into the effects of social media ranking algorithms, most are observational and have often yielded inconsistent findings (?). The few randomized experiments (?, ?, ?, ?) that do exist have been primarily conducted by the companies that design these algorithms. While these in-house studies provide valuable insights, they are typically infeasible for external researchers to replicate. This accentuates the need for more randomized experimentation, especially from outside of the companies.
Moreover, while it is essential to know whether these algorithms are causing adverse outcomes, to design better alternatives, we must also understand why they do so. Even existing randomized experiments that isolate the effect of the ranking algorithm seldom delve into identifying which specific elements of the algorithm contribute to such outcomes. In this paper, we aim to delve deeper, not only examining if current algorithms are exacerbating divisiveness but also exploring the extent to which this is driven by their prioritization of engagement (users’ revealed preferences) over a more reflective notion of users’ preferences.
To do so, we conducted a randomized controlled experiment measuring the impacts of Twitter’s [^1] Twitter rebranded to “X” in July 2023. Our study took place before the rebranding, so we refer to the platform as “Twitter”. engagement-based ranking algorithm. Our study was pre-registered, required no collaboration with the platform, and our data and code are available to the research community at http://github.com/smilli/twitter. Over two weeks in February 2023, we recruited a group of Twitter users (N=806) and collected the first ten tweets that the personalized algorithm would have shown them and the first ten tweets they would have seen under a reverse-chronological baseline. Using a within-subjects design, we showed users the tweets from both rankings in a randomized order and surveyed them about each one, allowing us to measure the causal impact of the engagement-based ranking on readers’ perceptions, relative to the reverse-chronological baseline.
To investigate the role of the engagement-based algorithm in shaping discourse and attitudes, we studied its effect on six key outcomes (Outcomes 1-6 in Table 1) related to the curation of emotional and political content and the emotions and political perceptions of readers. In particular, we measured how the algorithm affected the emotions, partisanship, and out-group animosity expressed in the content shown, as well as the emotions experienced by the reader, and the reader’s perception of their political in-group and out-group. Throughout our paper, we refer to these six outcomes as “sociopolitical outcomes”. As summarized in Table 1, there is little experimental evidence for many of these outcomes. Thus, our first line of inquiry can be summarized by the following research question:
@@ 6 Reader’s out-group perception  (?) Ours, (?, ?) Ours
    How does the engagement-based ranking algorithm impact the sociopolitical outcomes under consideration (Outcomes 1-6)?
Our second research question asked whether Twitter’s ranking algorithm, which optimizes primarily for engagement (users’ revealed preferences) also aligns with users’ explicitly stated preferences. Many hypothesize that an engagement-centric approach may inadvertently promote polarizing content by focusing on users’ immediate reactions rather than their reflective evaluations of the content (?, ?, ?, ?). Therefore, if this engagement-focused characteristic is a key factor in the amplification of divisiveness, it would be reasonable to expect that the algorithm might underperform in aligning with users’ stated preferences. However, we found almost no prior studies that provided evidence for how well ranking algorithms satisfied users’ stated preferences [^2] Jia et al. (2023) (?) was an exception. However, they did not evaluate a real-world recommender system’s impact on user satisfaction, like the Twitter or Facebook algorithm, but rather their own custom non-personalized ranking of Facebook posts based on public reaction counts. (Table 1), relative to simple baselines like a chronological or random feed.
@@
    Does the engagement-based ranking algorithm show more content that users say they want to see (Outcome 7)?
Finally, in exploratory analysis, we measured the impact of ranking based on users’ explicitly stated preferences rather than the traditional engagement-based approach. Aligning with a more reflective notion of value—focusing on users’ stated preferences—could potentially counter the negative effects commonly associated with engagement-based algorithms (?, ?, ?, ?). However, it is crucial to verify these theoretical benefits as it could be that users still prefer content that is divisive or hostile, even after reflection (?). We were unable to find any prior work measuring the effect of ranking by stated preferences on the sociopolitical outcomes we consider (Table 1). Thus, our last, exploratory inquiry aims to fill this gap and is summarized in the following question.
@@
    What impact does ranking by users’ stated preferences have on the sociopolitical outcomes considered (Outcomes 1-6)?
@@ 1 Results
From February 11 to February 27, 2023, we conducted our experiment with active Twitter users from the United States. We used CloudResearch Connect, an online crowd-working platform, to recruit participants. The study period was divided into four waves, and participants were allowed to take part in the study once during each wave. In total, 806 unique users participated in our study a total of 1730 times. Each time a user participated in our study, we collected two sets of tweets for them: (a) the first ten tweets they would have been shown by Twitter’s engagement-based algorithm and (b) the first ten tweets they would have seen if they were shown the latest tweets from the people they follow.
@@
Using a within-subjects design, we then surveyed users about tweets from both of their timelines in a randomized order, querying them about Outcomes 1-7 in Table 1.333If a tweet was a quote tweet or a reply to another tweet, then we surveyed the users about both the quote/quoted tweet and the reply/replied tweet. If a tweet was common to both the engagmeent-based and chronological timeline, we only surveyed the user about it once. For each tweet, we asked the user to assess whether it was about a political or social issue. If they marked the tweet as political, we asked them about the tweet’s ideological leaning, their own perception of their political in-group and out-group after reading the tweet, and whether the tweet author was expressing animosity towards their out-group. For all tweets (including non-political tweets), we asked users to assess the author’s emotions as well as the emotions the tweet made them feel, along four dimensions: happiness, anger, sadness, and anxiety. Finally, we also asked the reader whether they wanted to see tweets like it when they used Twitter.
Note that, for our main analysis, we rely on the reader’s judgments even for outcomes that are about the tweet itself (the tweet’s ideological leaning, the emotions expressed by the author, and whether the tweet expresses out-group animosity). In Section S4.6, we use GPT-4 (?), a large language model, to label these outcomes instead and find that all results are qualitatively the same.
@@ 1.1 Effects of engagement-based ranking
First, we state our findings on how Twitter’s engagement-based algorithm affects political outcomes, emotional outcomes, and the reader’s stated preference for the tweets shown. All tested outcomes and our analysis plan were pre-registered at https://osf.io/upw9a. Figure 1 shows a summary of the average treatment effect for each outcome. As specified in our pre-analysis plan, the average treatment effect is estimated through a difference in means (see Section S1.3), and two-sided $p$-values are estimated by paired permutation tests. In total, we tested 26 outcomes, and all results that are significant (at a $p$-value threshold of $0.05$) remain significant at a false discovery rate (FDR) of $0.01$. Thus, in expectation, none of our discoveries are false discoveries. The full table of standardized and unstandardized effect sizes, $p$-values, and FDR-adjusted $p$-values can be found in SM section S3.
@@
Figure 1: Average treatment effects for all outcomes. ATEs are shown with 95% Bootstrap confidence intervals (unadjusted for multiple testing). The effects of two different timelines are shown: (1) Twitter’s own engagement-based timeline, (2) our exploratory timeline that ranks based on users’ stated preferences. The effect sizes for both timelines are relative to the reverse-chronological timeline (the zero line). Average treatment effects are standardized using the standard deviation of outcomes in the chronological timeline (see Section S1.3 for details).
@@
Political outcomes (Outcomes 2, 3, 5, and 6). We defined the partisanship of a tweet as the absolute value of the reader-provided ideological leaning of the tweet (ranging from -2 = “Far left” to +2 = “Far right”). Relative to the reverse-chronological baseline, we found that the engagement-based algorithm amplified tweets that exhibited greater partisanship ($0.24$ SD, $p<0.001$) and expressed more out-group animosity ($0.24$ SD, $p<0.001$). Furthermore, tweets from the engagement-based algorithm made users feel significantly worse about their political out-group ($-0.17$ SD, $p<0.001$) and better about their in-group ($0.08$ SD, $p=0.0014$). These effects remained significant when considering left and right-leaning users specifically, except for in-group perception in left-leaning users, where we found no significant effect.
@@
Amplified emotionality (Outcomes 1 and 4). The algorithm significantly amplified tweets that expressed negative emotions—anger ($0.47$ SD, $p<0.001$), sadness ($0.22$ SD, $p<0.001$), and anxiety ($0.23$ SD, $p<0.001$). It also leads readers to feel more of all four emotions—anger ($0.27$ SD, $p<0.001$), sadness ($0.09$ SD, $p=0.003$), anxiety ($0.12$ SD, $p<0.001$), and happiness ($0.12$ SD, $p<0.001$). When considering only political tweets, we found that anger was by far the predominant emotion amplified by the algorithm, both in terms of the emotions expressed by authors ($0.75$ SD, $p<0.001$) and the emotions felt by readers ($0.37$ SD, $p<0.001$).
@@
User’s stated preference (Outcome 7). For each tweet, we also asked users whether they wanted to see tweets like it when they used Twitter. We found that overall, tweets shown by the algorithm are rated slightly higher ($0.06$ SD, $p=0.022$). Interestingly, however, the political tweets recommended by the personalized algorithm lead to significantly lower user value than the political tweets in the chronological timeline ($-0.18$ SD, $p=0.005$).

### v6

@@
Smitha Milli Cornell Tech, New York, NY, USA Corresponding author: smitha.milli@gmail.com Micah Carroll University of California, Berkeley, CA, USA Yike Wang University of Washington, Seattle, WA, USA
@@
University of California, Berkeley, CA, USA
Sebastian Zhao University of California, Berkeley, CA, USA Anca D. Dragan University of California, Berkeley, CA, USA
Classification: Social Sciences
Keywords: Social Media, Ranking Algorithms, Stated Preferences
@@
In a pre-registered algorithmic audit, we found that, relative to a reverse-chronological baseline, Twitter’s engagement-based ranking algorithm amplifies emotionally charged, out-group hostile content that users say makes them feel worse about their political out-group. Furthermore, we find that users do not prefer the political tweets selected by the algorithm, suggesting that the engagement-based algorithm underperforms in satisfying users’ stated preferences. Finally, we explore the implications of an alternative approach that ranks content based on users’ stated preferences and find a reduction in angry, partisan, and out-group hostile content, but also a potential reinforcement of pro-attitudinal content. The evidence underscores the necessity for a more nuanced approach to content ranking that balances engagement and users’ stated preferences.
1 Introduction
Social media ranking algorithms personalize content to individual users by predicting what they will engage with [19, 18]. For example, in April 2023, Twitter’s ranking algorithm was based on predicting whether a user would engage with a particular tweet using ten different types of engagement [50]. These predictions include behaviors like retweeting, replying, watching an embedded video, or lingering on a tweet for at least two minutes. Similarly, in December 2021, it was reported that TikTok’s algorithm predicted how long users would watch a video and whether they would like or comment on it [46]. We refer to these kinds of ranking algorithms as engagement-based ranking algorithms. These algorithms can be seen as optimizing for users’ revealed preferences, i.e., they interpret a user engaging with a piece of content as an indication that the user wants to see that piece of content [22].
There is concern that by focusing on maximizing user engagement—especially reactive or passive behaviors like lingering—these algorithms may amplify content that is attention-grabbing but is unaligned with the user’s reflective values or with broader societal values [51, 13, 34, 19, 1, 35]. For example, Brady et al. [13] suggest that, “content algorithms systematically exploit human social-learning biases because they are designed to optimize attentional capture and engagement time on the platform, and social-learning biases strongly predict what users will want to see.” They specifically claim that ranking algorithms tend to amplify what they term ‘PRIME’ content—prestigious, in-group, moral, and emotional content. While humans may have developed biases toward PRIME information for adaptive reasons, they suggest that in the context of modern social media, exploiting these biases can lead to heightened social misperception and conflict. Similarly, Agan et al. [1] suggest that since cognitive biases are more likely to be triggered during fast, intuitive thinking (“system 1”), algorithms trained on more passive or impulsive behaviors—like time spent watching a video or lingering on a tweet—are likely to replicate these biases.
The notion that algorithms exploit implicit human biases to drive user engagement suggests that, if users deliberately chose the content they consumed, it would differ from what the algorithm prioritizes. However, most existing research does not clearly disentangle the algorithm’s exploitation of automatic human biases from deliberate human choices. For example, many studies do not control for who the user follows. On most social media platforms, the content that is shown to a user is heavily shaped by their own explicit choices of who to follow. Many existing studies find a correlation between the engagement that a post receives and characteristics such as its emotional intensity [11, 12, 41]. However, it is possible that those posts received more engagement, not due to the algorithm prioritizing them, but because users chose to follow accounts that post more emotional content. Analogously, it was often hypothesized that YouTube’s recommendation algorithms led users down “rabbit holes” to politically extreme content [49, 43, 28, 47]. However, subsequent studies revealed that users often reach extreme channels through external links and subscriptions, suggesting that intentional human choice and demand might be an alternative explanation for the consumption of extreme content [17, 43, 31, 32].
Moreover, even controlling for who a user follows is not sufficient to disentangle deliberate choices from automatic biases. Most studies control for who a user follows by comparing posts selected by a ranking algorithm to those in a reverse-chronological baseline which consists of recent posts from the accounts a user follows. For example, in concurrent work, Bouchaud et al. [9] conducted such an audit of Twitter and found that the ranking algorithm selected more emotionally valent content compared to the reverse-chronological baseline. This implies that the ranking algorithm amplifies emotional content beyond what users’ deliberate choices of who to follow alone would generate, and therefore, provide stronger evidence of the algorithm exploiting human biases. However, such evidence would be far from conclusive. The algorithm might be surfacing more emotional content because, within the accounts a user follows, the user may actually prefer more emotional posts, even if they lack a direct mechanism to filter content at such a granular level. In other words, the algorithm could still be optimizing for what users would have deliberately selected on their own, given finer control over their content.
Disentangling deliberate human choice from algorithmic exploitation of automatic biases is important not only for scientific understanding but also for designing better systems. If as Brady et al. [13] suggest, the algorithmic exploitation of social biases leads to greater social misperception and conflict, then redesigning algorithms to avoid such exploitation becomes imperative. In fact, a growing body of research in computer science is working toward this goal. For example, Milli et al. [36] developed a method that optimizes for passive engagement only to the extent that they are consistent with and predictive of users’ stated content preferences. Agarwal et al. [2] utilize generative temporal point processes to create a so-called ‘System 2 recommender system’ which models when users return to the platform to assess whether recommendations align with their deeper, reflective preferences. Kleinberg et al. [35] provide a more theoretical analysis but suggest ways that platform can understand whether they are over-optimizing for System 1 while failing to provide reflective value. These approaches all operate under the assumption that algorithms exploit automatic biases and that correcting for this could provide greater value to individuals and society. To provide a stronger grounding and justification to these methods, it is crucial to empirically test this hypothesis, which is the aim of our work.
We executed a pre-registered algorithmic audit of Twitter [^1] Twitter rebranded to “X” in July 2023. Our study took place before the rebranding, so we refer to the platform as “Twitter”. that was designed to differentiate deliberate human choices from algorithmic exploitation of biases. Over two weeks in February 2023, we recruited a group of Twitter users (N=806) and collected (1) the first ten tweets the engagement-based personalized ranking algorithm would have shown them, and (2) the ten most recent tweets they would have seen from accounts that they follow, i.e., a reverse-chronological baseline. [^2] On Twitter, users can opt out of the engagement-based timeline and use the reverse-chronological baseline instead, making it an especially relevant baseline to test against. By comparing to the reverse-chronological baseline, we are able to understand the effects of the engagement-based timeline, beyond users’ own deliberate decisions of who to follow.
Moreover, beyond simply controlling for users’ decisions of who to follow, we also explicitly surveyed users regarding their preferences for the tweets shown in both their engagement-based and reverse-chronological timelines. In exploratory analysis, we used these survey responses to construct a counterfactual stated preference timeline consisting of the collected tweets that the user said they valued the most. The stated preference timeline simulates how a timeline might look if algorithms were more aligned with users’ stated preferences, i.e., preferences that are derived from what users say they want, rather than users’ revealed preferences which are based on which items they engage with. We then compared this stated preference timeline to that of the engagement-based and reverse-chronological timeline. We then compared this stated preference timeline with the engagement-based and reverse-chronological timelines, allowing us to assess whether the effects of the engagement-based algorithm were still more pronounced than what might occur if users could curate their content at the tweet level (rather than merely by whom they follow).
1.1 Hypotheses and research questions
We pre-registered five primary hypotheses and four secondary research questions. All hypotheses and research questions are stated as expectations about the effects of the engagement-based timeline relative to their reverse-chronological timeline. In exploratory analysis, we also analyzed the same outcomes for the stated preference timeline that we constructed using users’ explicit stated value for individual tweets.
Our first three hypotheses focused on the differences in the content selected by the engagement-based timeline and the reverse-chronological timeline. In particular, we hypothesized that, compared to the reverse-chronological timeline, the engagement-based timeline would show tweets (i) that were more emotional, (ii) had a stronger ideological leaning, and (iii) contained more expressions of animosity towards participants’ political out-group.
  1. Hypothesis 1.
    The engagement-based timeline will show tweets that are more emotional (along four dimensions of anger, sadness, anxiety and happiness), compared to the reverse-chronological timeline.
  1. Hypothesis 2.
    The engagement-based timeline will show tweets with a stronger ideological leaning, compared to the reverse-chronological timeline.
  1. Hypothesis 3.
    The engagement-based timeline will show tweets containing greater out-group animosity, compared to the reverse-chronological timeline.
These three hypotheses are consistent with both prior theoretical and empirical work. They are consistent with the theoretical framework put forth by Brady et al. [13], who suggest that social media ranking algorithms exploit humans’ social biases towards emotional, moralized, and in-group content. Moreover, previous observational studies have found that tweets with more emotion and out-group animosity tend to receive higher engagement [41, 11]. However, as previously noted, these studies do not account for users’ choices in whom they follow, making it challenging to assess the specific influence of the ranking algorithm in amplifying emotional and out-group antagonistic content. Our hypotheses posit that engagement-driven algorithms do indeed amplify such emotionally charged, politically divisive content beyond the impact of users’ own follow choices.
We also had two hypotheses about how participants would feel after reading the tweets in their engagement-based timeline. The first was that, by selecting tweets with greater ideological leaning and out-group animosity, we expected readers to feel worse about their political out-group after reading tweets from their engagement-based timeline.
  1. Hypothesis 4.
    Participants will feel worse about their political out-group after reading tweets from their engagement-based timeline, compared to the reverse-chronological timeline.
Our second reader-focused hypothesis was that participants would feel happier but less angry, sad, or anxious reading tweets in their engagement-based timeline. This contrasted with our expectation that the engagement-based timeline would increase the frequency of expressions of all four emotions (Hypothesis 1). The rationale for this difference is that platforms test ranking algorithms in A/B tests to identify those that optimize user retention [19]; thus, we assumed that fostering positive emotions would encourage users to return to the platform. (Ultimately this turns out to be the only hypothesis we did not find supporting evidence for — readers reported feeling higher levels of all four emotions on the engagement-based timeline.)
  1. Hypothesis 5.
    Participants will feel happier and less angry, anxious, or sad reading tweets in their personalized timeline, compared to the reverse-chronological timeline.
We also formulated four secondary pre-registered research questions. The first two explored the emotional impact of the engagement-based timeline on both the content it curates (as in Hypothesis 1) and the emotions experienced by readers (as in Hypothesis 4), but focused specifically on political tweets. While we expected that the engagement-based timeline would increase the expression of all four measured emotions—anger, happiness, sadness, and anxiety—across tweets overall (Hypothesis 1), we were uncertain about the effects when considering only political tweets. It is possible that political tweets might primarily amplify negative emotions rather than emotional content more broadly, with a similar pattern reflected in the emotions experienced by readers.
@@ 1.1 Hypotheses and research questions
    How angry, happy, sad, or anxious are the political tweets selected by the engagement-based timeline, compared to the reverse-chronological timeline?
@@
    How happy angry, happy, sad, or anxious do readers feel after reading political tweets selected by the engagement-based timeline, compared to the reverse-chronological timeline?
Our next question related to the reader’s change in perception of their political in-group. We hypothesized that the engagement-based timeline would make the reader feel worse about their out-group (Hypothesis 4) because of increased exposure to tweets expressing animosity to the out-group. However, we were uncertain whether, conversely, the reader’s perception of their in-group would improve when exposed to tweets from the engagement-based timeline, especially because previous research has found that users are more likely to share tweets that are negative towards their out-group than those that are positive towards their in-group [54].
@@
    How do readers’ perceptions of their political in-group change after reading tweets in their engagement-based timeline, compared to their reverse-chronological timeline?
Our final question investigated whether the engagement-based timeline aligns with users’ stated preferences—specifically, whether it shows them tweets they say they value. Users’ revealed preferences appear to favor the engagement-based timeline over the reverse-chronological one, as randomized experiments at Twitter have shown that it increases the amount of time users spend on the platform compared to the reverse-chronological timeline [37]. However, theories suggesting that ranking algorithms exploit social biases [13, 1] suggest that users’ revealed preferences, which are given in an automatic state, may significantly differ from their stated preferences, which are obtained in a more reflective way, e.g. through user surveys.
  1. Question 4.
    Do users see more tweets they (say they) value in the engagement-based timeline, compared to the reverse-chronological timeline?
@@ 2 Results
From February 11 to February 27, 2023, we conducted our study on CloudResearch Connect, an online crowd-working platform. The study period was broken into four waves [^3] The time periods for the waves (inclusive) were 02/11-02/14, 02/16-02/19, 02/21-02/23, and 02/25-02/27. and participants could complete the study once during each wave. Every day, we recruited up to 150 eligible participants who lived in the United States, were at least 18 years old, and used Google Chrome. Furthermore, participants were required to use Twitter at least a few times a week and follow at least 50 people on Twitter (both gauged through self-reports). To collect data, participants were directed to download a Chrome extension that we developed which scraped their Twitter homepage to collect the top tweets from their personalized, engagement-based timeline. While scraping, the Chrome extension added an overlay to the homepage that prevented the user from seeing the tweets during collection. At the same time that the engagement-based timeline was collected, we queried the Twitter API to get the top tweets from the reverse-chronological timeline. Only public tweets were collected and no promoted tweets (advertisements) were collected.
After collecting both sets of tweets, participants were directed to complete a survey on Qualtrics that asked questions about each of the top ten tweets from their engagement-based and reverse-chronological timeline. All tweets were displayed in a randomized order (thus, tweets from both timelines were typically interwoven rather than, for example, first showing all the engagement-based tweets and then all reverse-chronological tweets). If the same tweet was present in both the engagement-based and reverse-chronological timeline, then participants were only shown it once. If a tweet was a reply to another tweet, the user was shown both the replied tweet and the main tweet, and asked to answer the questions for both tweets. Similarly, if a tweet was a quote tweet, then the user was asked to answer the questions for both the quoted tweet and the main tweet.
@@
For each tweet, we asked the user to assess whether it was about a political or social issue. If they marked the tweet as political, we asked them about the tweet’s ideological leaning, their own perception of their political in-group and out-group after reading the tweet, and whether the tweet author was expressing out-group animosity. For all tweets (including non-political tweets), we asked users to assess the author’s emotions as well as the emotions the tweet made them feel, along four dimensions: happiness, anger, sadness, and anxiety. Finally, we also asked the reader for their stated preference about the tweet, i.e., whether they wanted to see tweets like it when they used Twitter. The full survey is provided in SM section S5.
For our analysis of outcomes that are about the tweet itself (the tweet’s ideological leaning, the emotions expressed by the author, and whether the tweet expresses out-group animosity), as opposed to the reader’s emotions, we measure these outcomes in two ways. First, as just described, we collect readers’ perceptions of these outcomes. Second, we ask the same questions to GPT-4 [39], a large language model (LLM). The main text presents the results based on readers’ judgments, and Section S4.6 includes results using GPT-generated labels. In the main text, we present the results based on readers’ judgments, while Section S4.6 contains results using GPT-provided labels. The findings indicate that both sets of results align qualitatively, providing reassurance given the limitations of each approach. Reader-given labels have the disadvantage that the reader’s own background, e.g. political leaning, education, etc, might cause them to systematically misinterpret certain outcomes such as the author’s emotions [14]. Conversely, while humans can grasp broader context beyond the tweet’s text, LLMs are limited to textual input and cannot process associated images or linked content. Moreover, large language models have their own biases in responses that may not be reflective of the human population [44, 20]. Nevertheless, the consistency between the two methods, each with their advantages and disadvantages, supports the robustness of our results.
@@ 2.1 Effects of engagement-based ranking
First, we state our findings on the effects of Twitter’s engagement-based algorithm. All tested outcomes and our analysis plan were pre-registered at https://osf.io/upw9a. Figure 1 shows a summary of the average treatment effect for each outcome. As specified in our pre-analysis plan, the average treatment effect (of the engagement-based algorithm, relative to the control of the reverse-chronological ranking) is estimated through a difference in means (see Section S1.3), and two-sided $p$-values are estimated by paired permutation tests. In total, we tested 26 outcomes, and all results that are significant (at a $p$-value threshold of $0.05$) remain significant at a false discovery rate (FDR) of $0.01$. Thus, in expectation, none of our discoveries are false discoveries. The full table of standardized and unstandardized effect sizes, $p$-values, and FDR-adjusted $p$-values can be found in SM section S3.
@@
Figure 1: Average treatment effects for all outcomes. ATEs are shown with 95% Bootstrap confidence intervals (unadjusted for multiple testing). The treatment effects of two different timelines are shown, relative to the reverse-chronological timeline: (1) Twitter’s own engagement-based timeline, (2) our exploratory timeline that ranks based on users’ stated preferences. The outcomes shown here are based on reader judgments, the analogous effects for GPT-4 based labels can be found in Section S4.6. The effect sizes for both timelines are relative to the reverse-chronological timeline (the zero line). Average treatment effects are standardized using the standard deviation of outcomes in the reverse-chronological timeline (see Section S1.3 for details).
@@
Political outcomes. We defined the partisanship of a tweet as the absolute value of the reader-provided ideological leaning of the tweet (ranging from -2 = “Far left” to +2 = “Far right”). Relative to the reverse-chronological baseline, we found that the engagement-based algorithm amplified tweets that exhibited greater partisanship ($0.24$ SD, $p<0.001$) and expressed more out-group animosity ($0.24$ SD, $p<0.001$). Furthermore, tweets from the engagement-based algorithm made users feel significantly worse about their political out-group ($-0.17$ SD, $p<0.001$) and better about their in-group ($0.08$ SD, $p=0.0014$). These effects remained significant when considering left and right-leaning users specifically, except for in-group perception in left-leaning users, where we found no significant effect.
@@
Amplified emotionality. The engagement-based algorithm significantly amplified tweets that expressed negative emotions—anger ($0.47$ SD, $p<0.001$), sadness ($0.22$ SD, $p<0.001$), and anxiety ($0.23$ SD, $p<0.001$). It also led readers to feel more of all four emotions—anger ($0.27$ SD, $p<0.001$), sadness ($0.09$ SD, $p=0.003$), anxiety ($0.12$ SD, $p<0.001$), and happiness ($0.12$ SD, $p<0.001$). When considering only political tweets, we found that anger was by far the predominant emotion amplified by the engagement-based algorithm, both in terms of the emotions expressed by authors ($0.75$ SD, $p<0.001$) and the emotions felt by readers ($0.37$ SD, $p<0.001$).
@@
User’s stated preference. For each tweet, we also asked users whether they wanted to see tweets like it when they used Twitter. We found that overall, tweets shown by the engagement-based algorithm are rated slightly higher ($0.06$ SD, $p=0.022$). Interestingly, however, the political tweets recommended by the engagement-based algorithm led to significantly lower user value than the political tweets in the reverse-chronological timeline ($-0.18$ SD, $p=0.005$).